he 2140-N block shown is in contact with 45° incline. The coefficient of static friction is 0.25. Compute the value of the horizontal force P necessary to just start the block up the incline. Write your answer in two decimal places.


1
Expert's answer
2021-03-02T18:09:38-0500

Let's apply the Newton's Second Law of Motion in projections on axis "x" and "y":


"Pcos\theta-F_{fr}-Wsin\theta=0,""N-Wcos\theta-Psin\theta=0."


Let's rewrite our equations:


"Pcos\theta-\mu_sN-Wsin\theta=0,""N=Wcos\theta+Psin\theta."


Let's substitute "N" into the previous equation and solve for "P":


"Pcos\theta-\mu_s(Wcos\theta+Psin\theta)-Wsin\theta=0,""P=\dfrac{W(sin\theta+\mu_scos\theta)}{cos\theta-\mu_ssin\theta},""P=\dfrac{2140\ N\cdot(sin45^{\circ}+0.25\cdot cos45^{\circ})}{cos45^{\circ}-0.25\cdot sin45^{\circ}}=3566.67\ N."